OpenStudy (anonymous):

f (x) is a differentiable function defined on the interval [0, 60]. x 0 15 30 45 60 f(x) -3 -1 2 6 2 Show that f '(c1) = 1/12 for some value x = c1 in the interval (0, 60). Justify your answer. I'm thinking that the Mean Value Theorem applies here... but I have no clue as to how to explain or verify this.. please anyone help? I've been trying to figure this out for hours, really...

OpenStudy (anonymous):

Yes, you should use the mean value theorem

OpenStudy (anonymous):

There is c between a and b so that f'(c)= (f(b)-f(a))/(b-a)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i mean i know whawt mean value theorem is... do you think that's all i need to say on this problem?

OpenStudy (anonymous):

(2-(-3)/(60-0) =5/60=1/12

OpenStudy (anonymous):

OH! i never thought of simplifying...

OpenStudy (anonymous):

a=0, b=60

OpenStudy (anonymous):

Wow. You are a genius! Thank you soooo much!!

OpenStudy (anonymous):

NP

OpenStudy (anonymous):

:D